# Relative Representations: Topological and Geometric Perspectives

### v1

@@ 1 Introduction and Related Work
Surprisingly, it has been argued that neural networks trained on diverse tasks, architectures, and domains, infer structurally similar representations [34, 25, 38] – a hypothesis sometimes referred to as ‘representational universality’ or ‘convergent learning’ [29]. For example, there is extensive empirical evidence that the representations extracted by neural networks are isometric – i.e., are related by an affine distance-preserving transformation – as the hyperparameters and initialization vary [35]. Even though partial theoretical explanations have been proposed – e.g., based on harmonic analysis [32, 11] or on philosophical principles [22] – these phenomena remain mostly mysterious. Nonetheless, they have motivated the introduction of techniques for zero-shot transfer and latent space communication, such as model stitching [28, 2, 12] – a method consisting in a trainable layer that connects latent representations of two different networks. On a similar note, relative representations [35, 7] involve a non-trainable isometry-invariant layer on top of the representation. Assuming the representational universality hypothesis, factoring out rigid transformations results in features that are independent of nuances in initialization, hyperparameters, and, to some extent, task and architecture.
@@ 2.1 Relative Representations
Relative representations enable zero-shot model stitching as follows. If $f_{1}=\gamma_{1}\circ\varphi_{1}$ and $f_{2}=\gamma_{2}\circ\varphi_{2}$ are two neural networks with the same architecture decomposed in a representation map $\varphi_{\bullet}$ and a head $\gamma_{\bullet}$, then one can replace $\gamma_{1}$ with $\gamma_{2}\circ T_{\text{rel}}^{\mathcal{A}}$ to exploit the representation inferred by $f_{1}$ in conjunction with the head of $f_{2}$, and vice versa. Assuming the representational universality hypothesis, this model stitching procedure extends to cross-domain setups, i.e., for two networks $f_{1}$ and $f_{2}$ that are trained over two distinct datasets from the same semantic domain (see Figure 1).
@@
In this case, to obtain coherent anchors, we fix $\mathcal{A}$ as an encoding via $\varphi_{1}$ of some data, and exploit a known correspondence between the datasets – based on domain knowledge – to obtain anchors $\mathcal{A}^{*}$ as encodings via $\varphi_{2}$ of the corresponding data. Lastly, for our purposes it will be convenient to resort to an end-to-end training setup, similarly to [27]. This consists in deploying pre-trained representation maps $\varphi_{1},\varphi_{2}$, and training the networks $\gamma_{1}\circ T_{\textnormal{rel}}^{\mathcal{A}}\circ\varphi_{1}$ and $\gamma_{2}\circ T_{\textnormal{rel}}^{\mathcal{A}^{*}}\circ\varphi_{2}$ with the weights of both representation maps unfrozen, using their corresponding domain-specific dataset.
@@ 3.1 Robust Relative Transformation
As a consequence, the robust relative transformation is invariant to the intertwiner groups induced by common activation functions (e.g. GELU, ReLU, sigmoid) – see discussion after Definition 2.2. Note that this invariance property differs from the one satisfied by the original version of the relative transformation (Definition 2.1). The latter is invariant to isotropic rescalings – or, equivalently, to diagonal matrices with equal diagonal entries – and linear isometries. Therefore, our version trades off invariance to isometries other than permutations with more general non-isotropic rescalings. We claim that this tradeoff is advantageous in high dimensions. An arbitrary orthogonal matrix can be approximated by a permutation one, with the error decreasing as the dimension grows [3]. Since latent spaces of contemporary deep learning models are typically high-dimensional, the robust relative transformation approximately exhibits, to an extent, invariance to arbitrary isometries, together with the added invariance to arbitrary non-isotropic rescalings. We empirically compare the robust version with the classical one as part of our experimental investigation – see Section 4.1.
@@ 4 Experiments
In this section, we provide an empirical investigation of our proposed improvements to relative representations. To this end, we train Transformer models on a natural language task. The task is split across languages, corresponding to different domains. The cross-domain stitching procedure enables to transfer between languages, i.e., it implements a form of zero-shot translation. Our experimental setup is similar to the original work [35].
@@
We perform a random subsampling of 1% of the Amazon Reviews dataset [24], where the task consists in predicting user ratings in a range from $1$ to $5$ stars. We consider such task in two languages: English (‘en’) and French (‘fr’). The anchor sets $\mathcal{A}$ and $\mathcal{A}^{*}$ (notation from Section 2.1) consist of translated texts. Specifically, we first select a number of anchors from the English dataset equal to the dimensionality of the latent space ($768$ dimensions), and then translate them to other languages using Google Translate. A Python code implementing our models and experiments is available at a public repository: https://github.com/AGarciaCast/TopoRobRelTrans.

### v2

@@ 1 Introduction and Related Work
Surprisingly, it has been argued that neural networks trained on diverse tasks, architectures, and domains, infer structurally similar representations [34, 25, 38] – a hypothesis sometimes referred to as ‘representational universality’ or ‘convergent learning’ [29]. For example, there is extensive empirical evidence that the representations extracted by neural networks are isometric (up to scale) – i.e., are related by an affine distance-preserving transformation – as the hyperparameters and initialization vary [35]. Even though partial theoretical explanations have been proposed – e.g., based on harmonic analysis [32, 11] or on philosophical principles [22] – these phenomena remain mostly mysterious. Nonetheless, they have motivated the introduction of techniques for zero-shot transfer and latent space communication, such as model stitching [28, 2, 12] – a method consisting in a trainable layer that connects latent representations of two different networks. On a similar note, relative representations [35, 7] involve a non-trainable isometry-invariant layer on top of the representation. Assuming the representational universality hypothesis, factoring out rigid transformations results in features that are independent of nuances in initialization, hyperparameters, and, to some extent, task and architecture.
@@ 2.1 Relative Representations
Relative representations enable zero-shot model stitching as follows. If $f_{1}=\gamma_{1}\circ\varphi_{1}$ and $f_{2}=\gamma_{2}\circ\varphi_{2}$ are two neural networks with the same architecture decomposed in a representation map $\varphi_{\bullet}$ and a head $\gamma_{\bullet}$, then one can replace $\gamma_{1}$ with $\gamma_{2}\circ T_{\text{rel}}^{\mathcal{A}}$ to exploit the representation inferred by $f_{1}$ in conjunction with the head of $f_{2}$, and vice versa.
@@
Assuming the representational universality hypothesis, this model stitching procedure extends to cross-domain setups, i.e., for two networks $f_{1}$ and $f_{2}$ that are trained over two distinct datasets from the same semantic domain (see Figure 1). In this case, to obtain coherent anchors, we fix $\mathcal{A}$ as an encoding via $\varphi_{1}$ of some data, and exploit a known correspondence between the datasets – based on domain knowledge – to obtain anchors $\mathcal{A}^{*}$ as encodings via $\varphi_{2}$ of the corresponding data. Lastly, for our purposes it will be convenient to resort to an end-to-end training setup, similarly to [27]. This consists in deploying pre-trained representation maps $\varphi_{1},\varphi_{2}$, and training the networks $\gamma_{1}\circ T_{\textnormal{rel}}^{\mathcal{A}}\circ\varphi_{1}$ and $\gamma_{2}\circ T_{\textnormal{rel}}^{\mathcal{A}^{*}}\circ\varphi_{2}$ with the weights of both representation maps unfrozen, using their corresponding domain-specific dataset.
@@ 3.1 Robust Relative Transformation
As a consequence, the robust relative transformation is invariant to the intertwiner groups induced by common activation functions (e.g. GELU, ReLU, sigmoid) – see discussion after Definition 2.2. Note that this invariance property differs from the one satisfied by the original version of the relative transformation (Definition 2.1). The latter is invariant to isotropic rescalings – or, equivalently, to diagonal matrices with equal diagonal entries – and linear isometries. Therefore, our version trades off invariance to isometries other than permutations with more general non-isotropic rescalings. We claim that this tradeoff is advantageous in high dimensions. According to [3], the action of an arbitrary orthogonal matrix on a normally distributed random vector $z\in\mathbb{R}^{m}$ can be approximated by a permutation of its coordinates, with a relative approximation error that diminishes as dimension increases at a rate of $\mathcal{O}((\ln m)/\sqrt{m})$. Consequently, despite the variability of permutation with respect to $z$, achieving full permutation invariance yields approximate invariance to arbitrary orthogonal transformations as $m\rightarrow\infty$. Since latent spaces of contemporary deep learning models are typically high-dimensional, the robust relative transformation approximately exhibits, to an extent, invariance to arbitrary isometries, together with the added invariance to arbitrary non-isotropic rescalings. We empirically compare the robust version with the classical one as part of our experimental investigation – see Section 4.1.
@@ 4 Experiments
In this section, we provide an empirical investigation of our proposed improvements to relative representations. To this end, we train Transformer models on a natural language task. The task is split across languages, corresponding to different domains. The cross-domain stitching procedure enables to transfer between languages, i.e., it implements a form of zero-shot translation.
@@
We perform a random subsampling of 1% of the Amazon Reviews dataset [24], where the task consists in predicting user ratings in a range from $1$ to $5$ stars. We consider such task in two languages: English (‘en’) and French (‘fr’). The anchor sets $\mathcal{A}$ and $\mathcal{A}^{*}$ (notation from Section 2.1) consist of translated texts. Specifically, we first select a number of anchors from the English dataset equal to the dimensionality of the latent space ($768$ dimensions), and then translate them to other languages using Google Translate. A Python code implementing our models and experiments is available at a public anonymized repository: https://tinyurl.com/TopoRelTrans.
@@
Our experimental setup is similar to the original work [35]. However, in order to test the models in a more challenging fine-tuning scenario, we train our models on fewer data (only 1% of the Amazon Reviews dataset). Moreover, we deploy a simpler linear head $\gamma$, instead of a deeper network. This naturally results in worse reported performance as compared to the original work.